OpenStudy (anonymous):

Consider the following reaction: 4NH3 (g) + 5O2 (g) 4NO (g) + 6H2O (g)

OpenStudy (anonymous):

O2 is the limiting reactant molecules NO = 4 x 35/5=28 molecules H2O = 6 x 35 /5= 42 molecules NH3 needed = 4 x 35 / 5 = 28 molecules NH3 in excess = 35 - 28 = 7 total molecules = 7 + 28 + 42 = 77